(1 point) Dumping and other material-handling operations near a landfill may result in contaminated particles being emitted into the surrounding air. To estimate such particulate emission, the following empirical formula can be used:

E(V,M)=k(0.0032)(V5)1.4(M2)−1.3E(V,M)=k(0.0032)(V5)1.4(M2)−1.3

where EE is the emission factor (in pounds of particles released into the air per ton of soil moved), VV is the mean speed of the wind (mph), MM is the moisture content of the material (given as a percentage), and kk is a constant that depends on the size of the particles.

(a) For a particle of a particular size, we find k=0.18k=0.18. Find E(10,18)E(10,18).

(c) The emission factor EE can be multiplied by the number of tons of material handled to obtain a measure of total emissions. Suppose 16 tons of the material in part (a) is handled. How many tons of a second kind of material with k=0.17k=0.17 and moisture content 12% must be handled to achieve the same level of total emissions if the wind velocity stays the same?

tons:
Answer #1

E(V,M) = k(0.0032)(V/5)^(1.4)*(M/2)^(-1.3)

k=0.18

(a)

E(10,18)= (0.18)(0.0032)(10/5)^(1.4)*(18/2)^(-1.3)

= 0.0000874

(b)

Equate both the equations

(16)(0.18)(0.0032)(2^1.4)(9^(-1.3)) = (T)(0.17)(0.0032)(2^1.4)((12/2)^(-1.3))

(16)(0.18)(9^(-1.3)) = T (0.17)(6^(-1.3)

On solving for T

T = 10 tons